OpenStudy (lena772):

he Ksp of Chromium (III) hydroxide is 1.60E-30. Calculate the molar solubility (M) 7.30e-16 (Tol: ± 1E-010) Answer Incorrect by more than 10%. concentration of chromium ion (M) (Tol: ± 1E-010) Answer Incorrect by more than 10%. concentration of hydroxide ion (M) 1.46e- (Tol: ± 1E-010) Answer Incorrect by more than 10%.

OpenStudy (lena772):

@matt101 i cant get this one for some reason, can you help me please

OpenStudy (matt101):

Let's do it step by step. First, write out the dissociation reaction for chromium (III) hydroxide.

OpenStudy (lena772):

Cr(OH)3--> 3OH-+ Cr+

OpenStudy (lena772):

is that correct?

OpenStudy (matt101):

Just about - keep in mind that Cr has a charge of +3! That won't affect how you tackle the rest of this particular question though. Now, based on that reaction, what is the general Ksp equation here?

OpenStudy (lena772):

Ksp = [Cr3+][OH-]^3?

OpenStudy (matt101):

Exactly right! Now let's find the molar solubility. What should the next step be?

OpenStudy (lena772):

1.60e-30=3x^2?

OpenStudy (lena772):

or is it 1.60e-30=x^4?

OpenStudy (lena772):

I believe I calculated it with 3x^2 and it was wrong

OpenStudy (lena772):

@matt101

OpenStudy (matt101):

So have a look at the format of the Ksp equation I sent you. Don't sub in any numbers just yet. Ksp=?

OpenStudy (lena772):

Ksp = [Cr3+] ^x + [OH-] ^y

OpenStudy (matt101):

And what are the values of x and y in this case?

OpenStudy (lena772):

x=1, y=3

OpenStudy (lena772):

I think

OpenStudy (lena772):

@matt101

OpenStudy (matt101):

Absolutely right! Now remember, our molar solubility is going to be the concentration of the ORIGINAL SOLID. How to the concentrations of each ion compare?

OpenStudy (lena772):

Molar solubility =1.60e-30, cation = 1.60e-30 anion= (1.60e-30)^3?

OpenStudy (matt101):

Close! First of all, your molar solubility is NOT the same as your Ksp. Your molar solubility is x, which is what we're solving for GIVEN the Ksp. That being said, the value of cation concentration here will be the same as the molar solubility, because each mole of solid contains one mole of cation. However, remember that the concentration of the anion is this case is THREE TIMES the molar solubility (3x), because every mole of the original solid contains THREE moles of the anion. This concentration, 3x, is THEN cubed as part of the Ksp equation. So your equation will be this: Ksp = [Cr3+]^1 [OH-]^3 1.60E-30 = (x)^1 (3x)^3 Now just solve this equation for x! Does this make more sense now?

OpenStudy (lena772):

Still not getting it right

OpenStudy (lena772):

molar solubility (M) 2.7E-8 (Tol: ± 1E-010) Answer Incorrect by more than 10%. concentration of chromium ion (M) 2.7e-8 (Tol: ± 1E-010) Answer Incorrect by more than 10%. concentration of hydroxide ion (M) 5.9e-23 (Tol: ± 1E-010) Answer Incorrect by more than 10%.

OpenStudy (matt101):

Ksp = [Cr3+]^1 [OH-]^3 1.60E-30 = (x)^1 (3x)^3 1.60E-30 = 27x^4 I simplified it by one more step. Can you solve it from there?